# Rotating twisted magnetosphere of magnetars: approximate analytical solutions

### v1

@@ Abstract
An approximate analytical solution for the rotating twisted magnetosphere of magnetars is presented. The poloidal flux is approximated by the self-similar twisted dipole field. The toroidal field is obtained by the minimum torque model. Under this approximation, it is found that: (1) The Y-point radius decreases with the increase of twist of the magnetic field. (2) The polar cap is larger for larger twist. (3) The particle outflow luminosity is larger for larger twist. (4) The maximum acceleration potential, pulse width of magnetar radio emission etc all increase with the twist. (5) For an untwisting magnetosphere, the physical properties general evolves towards that of the normal pulsars. The above findings are consistent with previous analytical and numerical results. The larger polar cap may corresponds to the hot spot during magnetar outburst. In general, a rotating twisted magnetosphere has larger open field line regions. The radio emission of magnetars and fast radio burst may both originates in the larger and evolving open field line regions of magnetars.
@@ 1 Introduction
The magnetosphere is the basic for pulsar multiwave emissions and spin-down torque. The change and evolution of the magnetosphere will also affect both the timing and pulse profile of pulsars (Kramer et al. 2006; Lyne et al. 2010; Desvignes et al. 2019). Traditionally, a static dipole field is often assumed in the study of pulsar electrodynamics and emission theories (Goldreich & Julian 1969; Ruderman & Sutherland 1975; Arons & Scharlemann 1979; Cheng et al. 1986; Qiao et al 2004). Up to now, a static dipole is also assumed in some studies of pulsar pulse profile and timing etc (Gangadhara 2004; Zhang et al. 2007; Cao et al. 2024; Roy et al. 2025).
@@
A self-consistent study of pulsar magnetosphere structure may be based on the “pulsar equation" (Mestel 1973; Michel 1973; Scharlemann & Wagoner 1973). Numerical solutions are obtained at the end of the 20th century (Contopoulos et al. 1999). There are many later developments (Gruzinov 2005; Timokhin 2006; Spitkovsky 2006; Philippov et al. 2015; Stefanou et al. 2023; see Contopoulos 2024a,b and references therein for an overview of previous works). The gamma-ray and radio emissions of pulsars may be model based this physical magnetosphere (Kalapotharakos et al. 2014; Philippov et al. 2020; see Philippov & Kramer 2022 for review).
@@
Compared with normal pulsars, magnetars may be neutron stars powered by their strong magnetic field (Duncan & Thompson 1992). The magnetic field in the case of magnetars carries free energy with them because the magnetic field may be twisted or have more complex geometry (Thompson et al. 2002; Pavan et al. 2009; Beloborodov 2009; Glampedakis et al. 2014; Akgun et al. 2016; Kojima et al. 2017; Tong 2019). The evolving magnetosphere of magnetars and/or crustal field is responsible for their bursts and outbursts (Coti Zelati et al. 2018; Tong & Huang 2020). Since magnetars typically have a rotation period about 10 seconds. Some modelings of magnetar magnetosphere assume that all the field lines are closed field lines and there is no open field lines in the case of magnetars. This will make the magnetosphere of magnetars very different from that of pulsars. However, observationally some high magnetic field pulsar also show magnetar activities (Gavriil et al. 2008; Gogus et al. 2016; Archibald et al. 2016). The radio emission of magnetars evolves from typical magnetars case to that of the normal pulsar case (Camilo et al. 2016; Huang et al. 2023). And it is a general belief that pulsars and magnetars belong to unified population of neutron stars (Rea et al. 2010; Vigano et al. 2013).
@@
The twisted magnetic field of magnetars may be described in a self-similar way: $B\propto r^{-(2+n)}$ (Thompson et al. 2002; Pavan et al. 2009; Tong 2019). Due to the twist of magnetic field, the field lines will inflate in the radial direction (Wolfson 1995; Thompson et al. 2002). Noting this point, by inducing the light cylinder radius, it is found that magnetars may have large polar caps (Tong 2019). Of course, the introduction of light cylinder radius is not a self-consistent treatment of the pulsar equation. Recently, numerical simulations of twisted magnetar magnetosphere are reported (Ntotikas et al. 2024, 2025). The authors found that (1)for a twisted magnetars, the Y-point radius can be much smaller than the light cylinder radius, (2) Consequently, the polar cap radius, spindown luminosity etc are much larger than the dipole case. Inspired by the numerical progresses, we try to provide an approximate analytical solution of the rotating twisted magnetosphere of magnetars.
@@
The existence of large polar cap and open field line region (Tong 2019; Ntotikas et al. 2024,2025) are not only related to the magnetosphere physics of magnetars and the unification of pulsars and magnetars. It may be also related to the physics of fast radio bursts. The observations of magnetar SGR 1935+2154 tell us magnetars may host fast radio bursts (CHIME/FRB Collaboration 2020; Bochenek et al. 2020). If there are large open field line region in the case of magnetar, then the fast radio burst can be generated in the open field line regions. Then its natural that they share similar physics to that of pulsars and magnetars, e.g., rotating vector model (Mckinven et al. 2025; Liu et al. 2025), polarization (Luo et al. 2020), radius to frequency mapping (Wang et al. 2019; Lyutikov 2020; Tong 2022) etc. Radio wave propagating in the open field line regions may also be easier to break out the magnetosphere (Beloborodov 2021; Qu et al. 2022).
@@
The structure of this paper is: In Section 2, the approximate analytical solution is provided. Its comparison with the numerical simulations is presented in Section 3. Its application to magnetar outburst, especially magnetar radio emissions, is discussed in Section 4. Section 5 is the conclusion.
@@ 2 Approximate analytical solutions
For an axisymmetric magnetosphere, the magnetic field may be describe by the poloidal flux function $\Psi$ (which is related to the poloidal magnetic field) and poloidal current $\Phi$ (which is related to the toroidal magnetic field), following the symbol system Chen & Zhang (2021), in spherical coordinate $(r,\theta,\phi)$, the magnetic field is:
@@
The force-free condition result in a partial differential for the flux function and current, the so-called “pulsar equation" (Scharlemann & Wagoner 1973):
@@
where $\Omega$ is the angular velocity of the compact star (assumed to be a constant), $\Phi^{\prime}\equiv d\Phi/d\Psi$. The first line in the above equation is named as the “non-rotational term", while the second line is named as the “rotational term" (Chen & Zhang 2021). Actually this pulsar equation finds its application in both pulsars and black holes [^1] In the black hole case, the angular velocity may not be a constant, e.g., for field lines anchered on the accretion disk. The pulsar equation will also contain additional terms which depends on $\Omega^{\prime}$. (See Narayan et al. 2007 and references therein.). For a nonrotating magnetar $\Omega=0$, as have been considered by some previous works, the pulsar equation reduces to the case of Grad-Shafranov equation (Wolfson 1995; Thompson et al. 2002; Pavan 2009; Tong 2019):
@@
The pulsar equation and the Grad-Shafranov equation can be solved numerically. At the same time, there also exist some approximate analytical solutions. These analytical solutions demonstrates the physics involved more easily.
When studying the analytical solution of pulsar equation for black hole jets and winds, it is found that (Narayan et al. 2007; Tchekhovskoy et al. 2008; Chen & Zhang 2021), (1) both the solutions $\Psi$ of the “non-rotation term" and the “rotation term" can be can be viewed as the approximate solution of the full pulsar equation. (2) The poloidal current (i.e., toroidal field) $\Phi$ is related to the flux function by the minimum torque model (Michel 1969; Narayan et al. 2007; Tchekhovskoy et al. 2008; Chen & Zhang 2021):
@@
where $0\leq n\leq 1$ is a parameter characterizing the twist of the magnetic field. Substituting eq.(5) into eq.(1), it can be seen that (Wolfson 1995), the magnetic field depends on the radius as $B\propto r^{-(2+n)}$. Therefore, the parameter $n=1$ corresponds to the dipole case, $n=0$ corresponds to the split monopole case. $0<n<1$ corresponds to the twisted dipole case. The function $f(\theta)\approx 1-\cos^{2}\theta=\sin^{2}\theta$ is the angular dependence of the flux function. As have been shown in previous works (Pavan et al. 2009; Tong 2019), the exact function of $f(\theta)$ is similar to the dipole case for small twist. This is especially true for $\theta=0$ (pole) or $\theta=\pi/2$ (equator), where the difference is even smaller. Similar things also happen in the black hole case (Tchekhovskoy et al. 2008), where the angular dependence can be approximated by split monopole case.
Equation (5) and (4) are the approximate analytical solutions to the pulsar equation, i.e., for a rotating twisted magnetosphere. A straightforward way to check this point is to substitute these two equation into the original pulsar equation. Considering previous experiences (Narayan et al. 2007; Tchekhovskoy et al. 2008; Chen & Zhang 2021), the approximation is rather good for small $\theta$ (i.e., the polar cap region). For other values of $\theta$, the same expression can only be viewed as approximations. The calculation is straightforward [^2] The blackhole case mainly focus on the case of flux increasing with radius: $\Psi=r^{\nu}f(\theta)$, where $0\leq\nu\leq 2$. Considering that $\nu$ is just a constant parameter, the calculations in the blackhole case are still valid at present. Except for this point (a different radial dependence), previous blackhole studies have already confirmed that eq.(5) and (4) are the approximate solutions to the pulsar equation..
The basic picture is that (similar to that in the numerical simulations, Ntotsikas et al. 2024, 2025): (1) In the closed field region, the poloidal field is describe by the twisted dipole field (eq.5). The toroidal field there is due to twist of the field, already fixed by the parameter $n$ for the self-similar case. (2) In the open field line region, the poloidal field is also described by the flux function (eq.5). The toroidal field is in the open field line region is due to rotation of the neutron star (eq.4). There are many assumptions and approximations in doing so. We will explore the consequence of this solutions and try to compare it with the up-to-date numerical solutions (Ntotsikas et al. 2024, 2025). In this way, we hope this approximate analytical solution can be justified.
@@ 2.1 The Y-point radius
Assuming the approximate analytical solution, from the definition of the magnetic field, the three components of the magnetic field are (for open field line regions):
@@
$B_{\phi}$ $=$ $\frac{1}{r\sin\theta}\Phi=-2\frac{r^{-(1+n)}}{R_{\rm lc}}\frac{1}{\sin\theta}f(\theta).$ (8)
The main quantities related to the function $f(\theta)$ is: $f(\theta=\pi/2)=1$ (conservation of flux), $f^{\prime}(\theta=\pi/2)=0$ symmetry about the equator. If needed, the approximation for $f(\theta)$ can be used: $f(\theta)\approx\sin^{2}\theta$. For a neutron star with surface polar magnetic field $B_{p}$, the correspond flux function should include a normalization factor $(1/2)B_{p}R^{2+n}$, where $R$ is the neutron star radius. Then at the surface of the neutron star, the flux function is: $\Psi(r=R)=(1/2)B_{p}R^{3}f(\theta)/R$. Neglecting a normalization factor $(1/2)B_{p}R^{3}$, the flux function at the surface is: $\Psi(r=R)=f(\theta)/R$. This expression is independent on the factor $n$, and is valid for both the dipole and twisted dipole case.
@@
Therefore, for a dipole field with $n=1$, the Y-point radius is $0.5$ times the traditional light cylinder radius. This may seem ridiculous at first sight. However, according to Contopoulos et al. (2024a,b), the Y-point radius may line at 0.80.90.8-0.9 the light cylinder radius. At present, we will neglect this effect and require that for a dipole field, the Y-point radius lines at the light cylinder radius. This will fix the numerical factor in the above express for the Y-point radius [^3] This equation can be easily extended to a form: $R_{Y}=ynR_{\rm lc}$, where $y$ may be in the range (0.80.9)(0.8-0.9) considering recent numerical results for a dipole field ($n=1$). The modifications to the following results are only quantitative. :
@@ 2.2 The polar cap
This can be compared with eq.(12) in Tong (2019), where the polar cap for a twisted magnetosphere is obtained by introducing the light cylinder radius: $\sin\theta_{\rm pc}\approx(R/R_{\rm lc})^{n/2}$. The present result (eq.(16)) differs from the that in Tong (2019) by a numerical factor in the denominator. Therefore, the previous conclusions in Tong (2019) are still valid (untwisting of a globally twisted magnetosphere due to physics in the open field line regions). The changes are only quantitative. However, Tong (2019) introduced the light cylinder radius as the starting point. Therefore, the Y-point radius can not be obtained self-consistently. While the present treatment obtained the Y-point radius (eq.12) using the approximation analytical solution to the pulsar equation. And the polar cap radius is obtained in a self-consistent way.

### v2

@@ Abstract
An approximate analytical solution for the rotating twisted magnetosphere of magnetars is presented. The poloidal flux is approximated by the self-similar twisted dipole field. The toroidal field is obtained by the minimum torque model. Under this approximation, it is found that: (1) The Y-point radius decreases with the increase of twist of the magnetic field. (2) The polar cap is larger for larger twist. (3) The particle outflow luminosity is larger for larger twist. (4) The maximum acceleration potential, pulse width of magnetar radio emission both increase with the twist. (5) For an untwisting magnetosphere, the physical properties evolve toward that of the normal pulsars. The above findings are consistent with previous analytical and numerical results. The larger polar cap may correspond to the hot spot during magnetar outburst. In general, a rotating twisted magnetosphere has larger open field line regions. The radio emission of magnetars and fast radio bursts may both originate in the larger and evolving open field line regions of magnetars.
@@ 1 Introduction
The magnetosphere is the basis for pulsar multiwave emissions and spin-down torque. The change and evolution of the magnetosphere will also affect both the timing and pulse profile of pulsars (Kramer et al. 2006; Lyne et al. 2010; Desvignes et al. 2019). Traditionally, a dipole field is often assumed in the study of pulsar electrodynamics and emission theories (Goldreich & Julian 1969; Ruderman & Sutherland 1975; Arons & Scharlemann 1979; Cheng et al. 1986; Qiao et al 2004). Up to now, a dipole field is assumed in some studies of pulsar pulse profile and timing (Gangadhara 2004; Zhang et al. 2007; Cao et al. 2024; Roy et al. 2025).
@@
A self-consistent study of pulsar magnetosphere structure may be based on the “pulsar equation" (Mestel 1973; Michel 1973; Scharlemann & Wagoner 1973). Numerical solutions are obtained at the end of the 20th century (Contopoulos et al. 1999). There are many later developments (Gruzinov 2005; Timokhin 2006; Spitkovsky 2006; Philippov et al. 2015; Stefanou et al. 2023; see Contopoulos 2024a,b and references therein for an overview of previous works). The gamma-ray and radio emissions of pulsars may be model based on this physical magnetosphere (Kalapotharakos et al. 2014; Philippov et al. 2020; see Philippov & Kramer 2022 for review).
@@
Compared with normal pulsars, magnetars may be neutron stars powered by their strong magnetic field (Duncan & Thompson 1992). The magnetic field in the case of magnetars carries free energy because the magnetic field may be twisted or have more complex geometry (Thompson et al. 2002; Pavan et al. 2009; Beloborodov 2009; Glampedakis et al. 2014; Akgun et al. 2016; Kojima et al. 2017; Tong 2019). The evolving magnetosphere of magnetars and/or crustal field is responsible for their bursts and outbursts (Coti Zelati et al. 2018; Tong & Huang 2020). Since magnetars typically have a rotation period about 10 seconds, some modelings of magnetar magnetosphere assume that all the field lines are closed field lines and there is no open field line in the case of magnetars. This will make the magnetosphere of magnetars very different from that of pulsars. However, observationally some high magnetic field pulsars also show magnetar activities (Gavriil et al. 2008; Gogus et al. 2016; Archibald et al. 2016). The radio emission properties of magnetars are different from that of normal pulsars. However, at the end of their radio activities, they evolve from typical magnetar case to that of the normal pulsar case (Camilo et al. 2016; Huang et al. 2023). And it is a general belief that pulsars and magnetars belong to a unified population of neutron stars (Rea et al. 2010; Vigano et al. 2013).
@@
The twisted magnetic field of magnetars may be described in a self-similar way: $B\propto r^{-(2+n)}$ (Thompson et al. 2002; Pavan et al. 2009; Tong 2019). Due to the twist of magnetic field, the field lines will inflate in the radial direction (Wolfson 1995; Thompson et al. 2002). Noting this point, by inducing the light cylinder radius, it is found that magnetars may have large polar caps (Tong 2019). Of course, the introduction of light cylinder radius is not a self-consistent treatment of the pulsar equation. Recently, numerical simulations of twisted magnetar magnetosphere are reported (Ntotikas et al. 2024, 2025). The authors found that for a twisted magnetosphere, the Y-point radius can be much smaller than the light cylinder radius. Consequently, the polar cap radius and spindown luminosity are much larger than the dipole case. Inspired by the numerical progress, we try to provide an approximate analytical solution of the rotating twisted magnetosphere of magnetars.
@@
The existence of large polar cap and open field line regions (Tong 2019; Ntotikas et al. 2024, 2025) are not only related to the magnetosphere physics of magnetars and the unification of pulsars and magnetars. It may also be related to the physics of fast radio bursts. The observations of magnetar SGR 1935+2154 tell us magnetars may host fast radio bursts (CHIME/FRB Collaboration 2020; Bochenek et al. 2020). If there are large open field line regions in the case of magnetars, then the fast radio bursts can be generated in the open field line regions. Then it is natural that they share similar physics to that of pulsars and magnetars, e.g., rotating vector model (Mckinven et al. 2025; Liu et al. 2025), polarization (Luo et al. 2020), radius to frequency mapping (Wang et al. 2019; Lyutikov 2020; Tong 2022). Radio wave propagating in the open field line regions may also be easier to break out the magnetosphere (Beloborodov 2021; Qu et al. 2022).
@@
The structure of this paper is: In Section 2, the approximate analytical solution is provided. Its comparison with the numerical simulations is presented in Section 3. Its application to magnetar outburst, especially magnetar radio emissions, is discussed in Section 4. The conclusion is presented in Section 5.
@@ 2 Approximate analytical solutions
For an axisymmetric magnetosphere, the magnetic field may be describe by the poloidal flux function $\Psi$ (which is related to the poloidal magnetic field) and poloidal current $\Phi$ (which is related to the toroidal magnetic field), following the symbol system in Chen & Zhang (2021), in spherical coordinate $(r,\theta,\phi)$, the magnetic field is:
@@
The force-free condition result in a partial differential equation for the flux and current, the so-called “pulsar equation" (Scharlemann & Wagoner 1973):
@@
where $\Omega$ is the angular velocity of the compact star (assumed to be a constant), $\Phi^{\prime}\equiv d\Phi/d\Psi$. The first line in the above equation is named as the “non-rotational term", while the second line is named as the “rotational term" [^1] In the closed field line regions, the toroidal field is due to twist of the magnetic field. While, in the open field line regions, the current and toroidal field are due to the rotation of the neutron star, see eq.(4). In the following, we are mainly interested in the open field line regions or the last open field line. Therefore, the $\Phi^{\prime}\Phi$ term is grouped into the “rotational term”. (Chen & Zhang 2021). Actually this pulsar equation finds its application in both pulsars and black holes [^2] In the black hole case, the angular velocity may not be a constant, e.g., for field lines anchered on the accretion disk. The pulsar equation will also contain additional terms which depends on $\Omega^{\prime}$. (See Narayan et al. 2007 and references therein.). For a nonrotating magnetar $\Omega=0$, as have been considered by some previous works, the pulsar equation reduces to the case of Grad-Shafranov equation (Wolfson 1995; Thompson et al. 2002; Pavan 2009; Tong 2019):
@@
The pulsar equation and the Grad-Shafranov equation can be solved numerically. At the same time, there also exist some approximate analytical solutions. These analytical solutions demonstrate the physics involved more easily.
When studying the analytical solution of pulsar equation for black hole jets and winds, it is found that (Narayan et al. 2007; Tchekhovskoy et al. 2008; Chen & Zhang 2021), (1) both the “non-rotation term" and the “rotation term" can be can be viewed as the approximate solution of the full pulsar equation. (2) The poloidal current is related to the flux function by the minimum torque model (Michel 1969; Narayan et al. 2007; Tchekhovskoy et al. 2008; Chen & Zhang 2021):
@@
where $0\leq n\leq 1$ is a parameter characterizing the twist of the magnetic field. Substituting eq.(5) into eq.(1), it can be seen that (Wolfson 1995), the magnetic field depends on the radius as $B\propto r^{-(2+n)}$. Therefore, the parameter $n=1$ corresponds to the dipole case, $n=0$ corresponds to the split monopole case. $0<n<1$ corresponds to the twisted dipole case. The function $f(\theta)\approx 1-\cos^{2}\theta=\sin^{2}\theta$ is the angular dependence of the flux function. As have been shown in previous works (Pavan et al. 2009; Tong 2019), the exact function of $f(\theta)$ is similar to the dipole case for small twist. This is especially true for $\theta=0$ (pole) or $\theta=\pi/2$ (equator), where the difference is smaller. Similar things also happen in the black hole case (Tchekhovskoy et al. 2008), where the angular dependence can be approximated by the split monopole case.
Equation (5) and (4) are the approximate analytical solution to the pulsar equation, i.e., for a rotating twisted magnetosphere. A straightforward way to check this point is to substitute these two equations into the original pulsar equation. Considering previous experiences (Narayan et al. 2007; Tchekhovskoy et al. 2008; Chen & Zhang 2021), the approximation is rather good for small $\theta$ (i.e., the polar cap region). For other values of $\theta$, the same expression can only be viewed as approximations. The calculation is straightforward [^3] The blackhole case mainly focus on the case of flux increasing with radius: $\Psi=r^{\nu}f(\theta)$, where $0\leq\nu\leq 2$. Considering that $\nu$ is just a constant parameter, the calculations in the blackhole case are still valid at present. Except for this point (a different radial dependence), previous blackhole studies have already confirmed that eq.(5) and (4) are the approximate solutions to the pulsar equation..
The basic picture is that (similar to that in the numerical simulations, Ntotsikas et al. 2024, 2025): (1) In the closed field line region, the poloidal field is describe by the twisted dipole field (eq.5). The toroidal field there is due to twist of the field, already fixed by the parameter $n$ for the self-similar case. (2) In the open field line regions, the poloidal field is also described by the flux function (eq.5). The toroidal field is in the open field line regions is due to rotation of the neutron star (eq.4). There are many assumptions and approximations in doing so. We will explore the consequences of this solution and try to compare it with the up-to-date numerical solutions (Ntotsikas et al. 2024, 2025). In this way, we hope this approximate analytical solution can be justified.
@@ 2.1 The Y-point radius
Assuming the approximate analytical solution, from the definition of the magnetic field, the three components of the magnetic fields are (for open field line regions):
@@
$B_{\phi}$ $=$ $\frac{1}{r\sin\theta}\Phi=-2r^{-(2+n)}\frac{r}{R_{\rm lc}}\frac{1}{\sin\theta}f(\theta).$ (8)
The main quantities related to the function $f(\theta)$ is: $f(\theta=\pi/2)=1$ (conservation of flux), $f^{\prime}(\theta=\pi/2)=0$ (symmetry about the equator). If needed, the approximation for $f(\theta)$ can be used: $f(\theta)\approx\sin^{2}\theta$. For a neutron star with surface polar magnetic field $B_{p}$, the correspond flux function should include a normalization factor $(1/2)B_{p}R^{2+n}$, where $R$ is the neutron star radius. Then at the surface of the neutron star, the flux function is: $\Psi(r=R)=(1/2)B_{p}R^{3}f(\theta)/R$. Neglecting a normalization factor $(1/2)B_{p}R^{3}$, the flux function at the surface is: $\Psi(r=R)=f(\theta)/R$. This expression is independent on the factor $n$, and is valid for both the dipole and twisted dipole case.
@@
Therefore, for a dipole field with $n=1$, the Y-point radius is $0.5$ times the traditional light cylinder radius. This may seem ridiculous at first sight. However, according to Contopoulos et al. (2024a,b), the Y-point radius may lie at 0.80.90.8-0.9 the light cylinder radius (see also particle-in-cell simulations, Philippov et al. 2015). In our opinion, there is an undetermined numerical factor in eq.(11). We require that for a dipole field, the Y-point radius equals the light cylinder radius. This will fix the numerical factor in the above express. And the final relation between the Y-point radius and the light cylinder radius is [^4] This equation can be easily extended to a form: $R_{Y}=ynR_{\rm lc}$, where $y$ may be in the range (0.80.9)(0.8-0.9) considering recent numerical results for a dipole field ($n=1$). The modifications to the following results are only quantitative.:
@@ 2.2 The polar cap
This can be compared with eq.(12) in Tong (2019), where the polar cap for a twisted magnetosphere is obtained by introducing the light cylinder radius: $\sin\theta_{\rm pc}\approx(R/R_{\rm lc})^{n/2}$. The present result (eq.(16)) differs from the that in Tong (2019) by a numerical factor in the denominator. Therefore, the previous conclusions in Tong (2019) are still valid (untwisting of a globally twisted magnetosphere due to physics in the open field line regions). The changes are only quantitative. However, Tong (2019) introduced the light cylinder radius as the starting point. Therefore, the Y-point radius can not be obtained self-consistently. While here the Y-point radius (eq.12) is obtained using the approximate analytical solution to the pulsar equation. And the polar cap radius is obtained in a self-consistent way.